# Imitation Learning for Combinatorial Optimisation under Uncertainty

### v1

@@ 3.1 Learning Paradigms
We describe two learning paradigms within IL: behavioural cloning and inverse reinforcement learning, followed by practical challenges and common remedies.
@@ 3.1.1 Behavioural Cloning (BC)
Behavioural cloning maps states, observations, or contexts to actions using expert demonstrations. The objective is to learn parameters $\theta$ by solving
@@
Equation˜5 minimises the one-step deviation error along expert trajectories. Within CO, two prominent approaches are common.
  • Parameter Prediction Task (PPT) - An ML model predicts parameters that define a CO problem, which is then solved by an external optimiser. Regret (also referred to as task loss in [donti2017task]) is typically defined as the difference between the optimal objective value with the knowledge of $\xi$ and the objective value realised by the prescriptive decision. Given state $x$ and a predictive model $\varphi_{\theta}$, we define the regret loss following [sadana2025survey] as
    $\mathcal{L}_{R}(x,a^{*},\theta)=C(x,a,\xi)-C(x,a^{*},\xi),$ (6)
    where $a=g(\varphi_{\theta}(x))$, $C$ denotes the total cost, and $g$ represents the optimisation solver. The loss in Equation˜6 measures the suboptimality of the induced decision $a$ relative to the expert action $a^{*}$ under realisation $\xi$. Training typically requires embedding a CO solver within the learning loop, often through a differentiable or problem-specific surrogate. The first part of the right-hand side of (6) ensures that the computed gap always considers the difference between the best achievable $a$ under $\theta$ and the optimal decision $a^{*}$. This gap always remains positive unless $a$ matches $a^{*}$, where the gap is zero. Hence, optimising this loss allows us to adjust $\theta$, to ensure that the CO-layer solution converges to the optimal solution. PPT is also referred to as the regret minimisation task (RMT) [sadana2025survey].
  • Action Imitation Task (AIT) - In this setting, a model $\pi_{\theta}$ directly outputs an action $a$ that approximates the expert action $a^{*}$, thereby reducing dependence on an external solver at inference time. A common imitation loss is
    $\mathcal{L}_{I}(x,a^{*},\theta)=\ell(a,a^{*}),$ (7)
    where $a=\pi_{\theta}(x)$ and $\ell(\cdot,\cdot)$ may be mean squared error or binary cross-entropy. When the regret objective is used as $\ell$, the approach aligns closely with PPT. AIT is also referred to as learning to optimise (LtO) [kotary2024learning].
@@
Figure 2: AIT or LtO: Loss = $\ell({a},{a^{*}})$
Comparing PPT and AIT
Under the PPT paradigm (Figure˜1), the ML model predicts parameters that define a combinatorial optimisation problem, which is then solved by an external optimiser. In contrast, the AIT framework (Figure˜2) eliminates this two-stage inference process by training the ML model to directly output a decision or action, thereby avoiding the need for a downstream optimisation solver at inference time. It is important to note, however, that an optimisation solver remains necessary during data generation to compute the expert actions $a^{*}$.
Imitation learning with AIT models in CO settings presents several challenges, primarily due to large and complex decision spaces and the presence of ambiguous demonstrations. Without explicitly structured output spaces, standard models may generate infeasible or inconsistent actions. Feasibility can be restored through post-processing techniques such as projection or heuristic correction, although this may result in a modest increase in regret [kotary2024learning]. The PPT approach avoids infeasibility by construction, as feasibility is enforced by the optimisation solver following parameter prediction [greif2024combinatorial, baty2024combinatorial]. This advantage comes at the cost of tighter coupling between the learning model and the solver, since the loss depends on the optimisation output and often requires differentiation through the solver. Such differentiation is typically complex and problem-specific, although gradient-based approaches remain popular due to their compatibility with neural network architectures [mandi2024decision, sadana2025survey]. Both PPT and AIT offer promising avenues for decision making in CO problems under uncertainty, yet practical understanding of their relative strengths remains limited.
@@
Given an offline training dataset $\mathcal{D}=\{(x_{i},a_{i}^{*})\}_{i=1}^{N}$ generated by an expert policy, that is $(x_{i},a_{i}^{*})\sim\pi^{*}$, behavioural cloning seeks to solve Equation˜5. This formulation gives rise to a distribution shift, or distribution mismatch, a phenomenon well documented in the supervised learning literature for sequential decision-making problems. During deployment, deviations from expert behaviour may lead the learned policy to encounter states that were absent from the training data. Such deviations can compound over time, causing the policy to drift into unfamiliar regions of the state space where reliable recovery is difficult.
@@
A standard methodology for addressing this mismatch is dataset aggregation (DAgger), introduced by [ross2011reduction]. DAgger iteratively generates trajectories using the current learned policy and relabels the visited states with actions provided by the expert. This process aligns the training distribution with the deployment distribution. However, in multi-stage CO settings, DAgger can be computationally expensive, as expert labelling requires repeatedly solving optimisation problems at many decision epochs. As a result, DAgger involves intermittent but costly expert feedback throughout the training process.
@@ 3.1.2 Inverse Reinforcement Learning (IRL)
Inverse reinforcement learning (IRL), first proposed in [russell1998learning], seeks to recover the underlying reward function from expert demonstrations rather than learning a direct mapping from states to actions. Consistent with the IRL literature, we use reward-based notation, in contrast to the cost-based formulation adopted for behavioural cloning. Once the reward function is inferred, a policy can be learned by solving the resulting reinforcement learning problem.
The objective is to recover a reward function parameterised by $\theta$ such that the induced optimal policy matches the expert policy $\pi^{*}$, that is,
@@
Given an expert dataset $\mathcal{D}$, the reward parameters are obtained by solving
@@
Reinforcement learning agents improve their behaviour through repeated interaction with the environment, allowing them to account for the long-term consequences of actions and recover more effectively from mistakes. As a result, IRL methods, which rely on RL for policy optimisation, are generally less sensitive to covariate shift than behavioural cloning. However, IRL is computationally demanding, as it requires repeated interaction with the environment and alternates between reward estimation and policy learning. IRL is also sample inefficient and inherently ill posed, since multiple reward functions can explain the same expert behaviour, making unique recovery of the true reward difficult.
Several classes of IRL methods have been proposed, differing mainly in how the objective function $\mathcal{J}$ is defined and optimised. Prominent approaches include maximum margin methods, which maximise the separation between expert and non-expert trajectories; maximum entropy IRL, which matches expert behaviour while encouraging stochasticity; Bayesian IRL, which infers a distribution over reward functions; and generative adversarial imitation learning (GAIL), which frames learning as a game between a policy and a discriminator. Comprehensive surveys of these methods are provided in [zare2024survey] and [zheng2022imitation].
IRL is particularly well-suited to settings where expert demonstrations are limited and accurate modelling of long-term behaviour is essential. A key advantage is that once the reward function is learned, policies can be re-optimised if system dynamics change, without requiring additional demonstrations. By contrast, behavioural cloning is more appropriate when expert data are abundant and cover the relevant state–action space, and when small deviations from expert behaviour do not lead to severe performance degradation.
@@ 3.2 Learning from Noisy or Suboptimal Experts
Many IL tasks rely on expert demonstrations, yet high-quality and fully optimal data are often difficult to obtain in practice. Human experts may make mistakes, and computing optimal solutions can be prohibitively expensive. Consequently, most real-world IL applications rely on imperfect but readily available demonstrations. In combinatorial optimisation (CO), human experts are typically unsuitable due to problem complexity, and optimisation solvers are used instead. However, these solvers are computationally demanding, particularly for sequential decision problems under uncertainty [powell2022reinforcement]. To mitigate this cost, demonstrations are often generated using static approximations or simplified scenarios. For example, [pham2023prediction, baty2024combinatorial] solve static CO problems offline to produce training targets. Although these solutions are not fully optimal, they are informative and substantially more tractable.
@@
In such settings, suboptimal demonstrations are frequently treated as if they were optimal, and classical IL methods are applied directly [pham2023prediction, baty2024combinatorial]. However, given the $\mathcal{NP}$-hard nature of many CO problems, even static instances often require approximations. Solvers are commonly run with time limits, heuristics, or prescribed optimality gaps, all of which introduce additional noise into the demonstrations. Despite this prevalence, there is limited understanding of how such imperfections affect IL performance in CO settings, and systematic comparative studies remain scarce.
@@
Several recent works explicitly consider learning under noisy supervision and adapt IL algorithms accordingly. [xu2022discriminator, kim2021demodice] propose offline IL frameworks that leverage supplementary datasets generated by suboptimal policies to enhance learning. [sekhari2024selective] introduces an IL algorithm that achieves low regret while reducing expert queries, even in the presence of noisy demonstrations. [wang2023imitation] presents a two-stage approach that first purifies noisy demonstrations using a diffusion process before applying imitation learning. Additional studies on learning from imperfect demonstrations include [wu2019imitation, sasaki2020behavioral, wang2021learning], as surveyed in [zare2024survey].
@@
Another challenge arises when multiple labels are available for the same state, often generated by different noisy experts. For instance, drivers may follow different policies under identical conditions, resulting in conflicting demonstrations. While such diversity can enrich the learning signal, it necessitates aggregation or filtering mechanisms to prevent degradation in policy performance. Early work by [raykar2009supervised, raykar2010learning] examines this issue in pre-labelled datasets. More recently, [sun2023mega] propose MEGA-DAgger, an interactive learning framework designed for multiple imperfect experts, which filters unsafe demonstrations and evaluates experts using scenario-specific criteria. [sekhari2024selective] further extend their approach to multi-expert settings in which each expert specialises in a distinct region of the state space.
@@ 3.3 Applications of IL in SDPs
Recent research on applying machine learning to SDPs in combinatorial optimisation largely follows two complementary directions. The first direction focuses on ML-augmented optimisation, where learning methods are used to enhance the efficiency of classical optimisation algorithms such as branch-and-bound or column generation. The second direction, referred to as end-to-end learning (E2EL), integrates combinatorial solvers or optimisation layers directly within learning architectures. This classification follows the taxonomy proposed by [kotary2021end]. While both IL and RL are employed across these approaches, the discussion here emphasises recent work that applies IL to decision making within these two categories.
@@ 3.3.1 ML-Augmented Optimisation
ML-augmented optimisation incorporates machine learning models into the decision-making processes of optimisation algorithms. In this paradigm, ML may act as a preprocessing tool to configure or initialise an algorithm, or as a component invoked repeatedly during execution to guide key decisions [bengio2021machine]. This approach has been explored across a wide range of problem domains. Some methods focus on predicting discrete actions, such as branching or node selection decisions, while others predict scores or surrogate quantities associated with these actions.
@@
    Cutting Planes. Cut generation is a central technique in mixed-integer programming, aimed at tightening relaxations without excluding feasible integer solutions. Several studies use supervised learning to decide which cuts to add, including [deza2023machine, ding2020accelerating, huang2022learning, paulus2022learning].
@@
    Column Generation and Branch-and-Price. Column generation solves large-scale linear programmes by generating variables on demand. When combined with Branch-and-Bound, this yields the Branch-and-Price framework [barnhart1998branch]. Learning-based methods in this setting focus on selecting promising columns or designing effective branching strategies to improve speed. Representative works include [furian2021machine, morabit2021machine, pereira2022learning, sun2022learning, shen2022enhancing].
@@
    Large Neighbourhood Search. Large Neighbourhood Search (LNS) improves solutions by iteratively destroying and repairing partial solutions. Machine learning has been used to generate high-quality initial solutions and to adaptively select neighbourhoods [hizir2025large, huang2023searching, liu2022learning, mossina2019multi, Nair2020SolvingMI, song2020general, sonnerat2021learning].
@@
    Search Space Reduction. Another line of work aims to directly reduce the search space by eliminating decisions unlikely to appear in an optimal solution. For example, in the travelling salesman problem, ML models can predict irrelevant arcs for pruning. Studies following this approach include [fischetti2019machine, fitzpatrick2021learning, grassia2019learning, lauri2019fine, larocca2024combining, morabit2022machine, spieckermann2025reduce, sun2021generalization, tahir2021improved, tayebi2022learning, vaclavik2018accelerating].
@@
    Constraint Programming. Constraint Programming (CP) solves combinatorial problems through systematic constraint propagation and search. Several works leverage supervised learning to guide CP solvers by predicting constraint satisfaction or search decisions, including [selsam2018learning, sun2024enhancing, xu2018towards].
@@
    Algorithm Configuration. Machine learning is also used to decide whether specific algorithmic components should be applied to a given instance. Examples include predicting when to apply Dantzig–Wolfe decomposition [kruber2017learning], when to use scaling techniques [berthold2021learning], whether to apply cutting planes beyond the root node [berthold2206learning], and whether to linearise the objective in mixed-integer quadratic programmes [bonami2018learning]. In all cases, algorithm runs are used to generate labelled training data.

### v2

@@ 3.1 Learning Paradigms
We describe two principal learning paradigms within IL: behavioural cloning and inverse RL, followed by a discussion of practical challenges and common mitigation strategies.
@@ 3.1.1 Behavioural Cloning (BC)
Behavioural cloning [pomerleau1988alvinn] formulates imitation as a supervised learning problem that maps states or observations directly to expert actions. The objective is to learn the parameters $\theta$ of a policy $\pi_{\theta}$ by solving
@@
Equation (1) minimises the one-step deviation between the learner and the expert along demonstrated trajectories. In the context of CO, two behavioural cloning formulations are commonly employed.
  • Parameter Prediction Task (PPT) - In this approach, an ML model predicts parameters that define a CO problem, which is subsequently solved by an external optimiser. The idea is to predict parameters that produce solutions close to the true ones in an imitation learning setting, where the expert knows the true parameters. The training objective is typically expressed in terms of regret (also referred to as task loss in [donti2017task]) when the predicted parameters appear in the objective of the problem. Regret is defined as the difference between the objective value achieved by the predicted decision and that of the expert decision under a realisation of uncertainty $\xi$. Following [sadana2025survey], the regret loss is given by
    $\mathcal{L}_{R}(x,a^{*},\pi_{\theta})=C(x,\pi_{\theta}(x),\xi)-C(x,a^{*},\xi),$ (2)
    where $C$ denotes the cost function. The loss in (2) is non-negative and vanishes only when the predicted decision coincides with the expert action. This paradigm is also referred to as the regret minimisation task (RMT) [sadana2025survey]. When predicted parameters appear in the constraints of the problem, regret alone is not sufficient, and one also needs to minimise the likelihood of infeasible outcomes. In such cases, the loss function involves a combination of a penalty for having infeasibility and losing optimality [mandi2025feasibility].
  • Action Imitation Task (AIT) - In contrast, AIT trains a model $\pi_{\theta}$ to directly output an action that approximates the expert decision, thereby eliminating the need for an optimisation solver at inference time. A common imitation loss is
    $\mathcal{L}_{I}(x,a^{*},\pi_{\theta})=\ell(\pi_{\theta}(x),a^{*}),$ (3)
    where $a=\pi_{\theta}(x)$ and $\ell(\cdot,\cdot)$ may be mean squared error or binary cross-entropy. When regret is used as the distance metric, AIT closely aligns with PPT. This approach is also known as learning to optimise (LtO) [kotary2024learning].
Comparing PPT and AIT
Under the PPT paradigm, the ML model predicts parameters of an easier CO problem that is subsequently solved by an external optimiser. By contrast, AIT eliminates this two-stage inference pipeline by training the model to directly output decisions, avoiding solver calls at deployment time. Nevertheless, an optimisation solver remains necessary during data generation to compute expert actions $a^{*}$. As a result, both paradigms rely on expert optimisation, but differ in when optimisation is used.
Figure 1: PPT
@@
IL with AIT models in CO settings presents several challenges, primarily due to large and complex decision spaces and the presence of ambiguous demonstrations. Without explicitly structured output spaces, standard models may generate infeasible or inconsistent actions. Feasibility can be restored through post-processing techniques such as projection or heuristic correction, although this may result in a modest increase in regret [kotary2024learning].
The PPT approach can mitigate infeasibility when the predicted parameters appear only in the objective function. However, when predicted parameters enter the constraints, the resulting solution may violate feasibility with respect to the constraints defined by the true parameters. Similar to AIT, [hu2023predict+] generate an initial solution using predicted parameters and subsequently apply corrective actions, if necessary, at an additional cost.
A key challenge in PPT lies in differentiating the loss function, as it depends on the optimisation outcome. Common strategies to address this include regularisation, perturbation techniques, and the use of surrogate loss functions [mandi2024decision, sadana2025survey]. Both PPT and AIT therefore offer promising approaches for decision making in CO under uncertainty; however, a systematic understanding of their relative strengths and limitations remains limited.
@@
- Given an offline training dataset $\mathcal{D}=\{(x_{i},a_{i}^{*})\}_{i=1}^{N}$ generated by an expert policy, that is $(x_{i},a_{i}^{*})\sim\pi^{*}$, behavioural cloning seeks to solve Equation˜1. This formulation gives rise to a distribution shift, or distribution mismatch, a phenomenon well documented [ross2011reduction] in the supervised learning literature for sequential decision-making problems. During deployment, deviations from expert behaviour may lead the learned policy to encounter states that were absent from the training data. Such deviations can compound over time, causing the policy to drift into unfamiliar regions of the state space where reliable recovery is difficult.
@@
A standard methodology for addressing this mismatch is dataset aggregation (DAgger) Algorithm˜1, introduced by [ross2011reduction]. DAgger iteratively generates trajectories using the current learned policy $\pi_{i}$ and labels the visited states with actions provided by the expert (steps 4 and 5). The action to be executed at any epoch is given by a function called the decision rule ($DR(\cdot)$) in step 7. An example of the Vanilla DR is given in Algorithm˜2. Vanilla DAgger executes the expert action with a decaying probability and the learner action otherwise, gradually shifting control from the expert to the learned policy over training iterations. At the end of the episode, a new policy denoted by $\hat{\pi}_{i+1}$ is trained on the updated dataset. By repeatedly collecting expert labels on states visited by the learner, DAgger aligns the training distribution with the state distribution induced at deployment. However, a key limitation in CO settings is the high computational cost of step 5, as expert labelling typically involves repeatedly solving CO problems at each decision epoch. We revisit the DAgger algorithm in Section˜5.
1
2Function DAgger():
3   Initial policy $\hat{\pi}_{0}$; expert $\pi^{*}$; iterations II; horizon $K$; learning routine $\textsc{Train}(\cdot)$; $\mathcal{D}\leftarrow\emptyset$ for $i\leftarrow 0$ to I1I-1 do
4      Observe initial state $x_{0}$;
    /* Collect a K-step trajectory */
5    for $k\leftarrow 0$ to K1K-1 do
       $a^{*}_{k}\leftarrow\pi^{*}(x_{k})$     // Query expert action
6
       $\mathcal{D}\leftarrow\mathcal{D}\cup\{(x_{k},a^{*}_{k})\}$  // Append demonstration
7
       $a_{k}\leftarrow DR(x_{k},a^{*}_{k},\hat{\pi}_{i})$  // Select action using decision rule
8
       $x_{k+1}\leftarrow X^{M}(x_{k},a_{k},\xi_{k})$// Environment transition
9
10      end for
11
    /* Retrain policy on aggregated dataset */
12    $\hat{\pi}_{i+1}\leftarrow\textsc{Train}(\mathcal{D})$;
13
14   end for
15
16 end
Algorithm 1 DAgger (Dataset Aggregation)
1Function VanillaDAgger($x_{k},{a}_{k},\hat{\pi}_{i}$):
2
3 $\beta_{i}\leftarrow\lambda^{i-1}\beta_{0}$;
4 $z\sim\text{Uniform}(0,1)$;
5
6 if $z\leq\beta_{i}$ then
7    return ${a}_{k}$;
8
9   end if
10 return $\hat{\pi}_{i}(x_{k})$;
11
12 end
Algorithm 2 VANILLA DAgger Decision Rule
@@ 3.1.2 Inverse Reinforcement Learning (IRL)
IRL, first introduced by [russell1998learning], aims to recover the underlying reward function that explains expert behaviour, rather than directly learning a mapping from states to actions. Following standard practice in the IRL literature, we adopt reward-based notation, in contrast to the cost-based formulation used for behavioural cloning. Once the reward function is inferred, a policy can be obtained by solving the resulting RL problem.
Formally, IRL seeks a reward function parameterised by $\theta$ such that the optimal policy induced by this reward coincides with the expert policy $\pi^{*}$:
@@
Given an expert dataset $\mathcal{D}$, the reward parameters are estimated by solving
@@
Because policy optimisation in IRL is typically performed using RL, IRL methods naturally account for long-term consequences of actions and are therefore less sensitive to covariate shift than behavioural cloning. However, this robustness comes at a high computational cost. IRL is generally sample inefficient, requires repeated interaction with the environment, and alternates between reward estimation and policy optimisation. Moreover, IRL is inherently ill-posed, since multiple reward functions may induce the same expert behaviour, making the unique recovery of the true reward function impossible in general.
A variety of IRL methods have been proposed, differing primarily in how the objective function $\mathcal{J}$ is defined and optimised. Notable approaches include maximum margin methods, which separate expert and non-expert trajectories; maximum entropy IRL, which matches expert behaviour while encouraging stochasticity; Bayesian IRL, which infers a distribution over reward functions; and generative adversarial imitation learning (GAIL), which frames imitation as an adversarial learning problem. Comprehensive surveys of these methods are provided in [zare2024survey] and [zheng2022imitation].
In practice, IRL is most suitable when expert demonstrations are limited and accurate modelling of long-term behaviour is essential. A key advantage is that once a reward function is learned, policies can be re-optimised under changing system dynamics without additional demonstrations. By contrast, behavioural cloning is preferable when expert data are abundant and sufficiently cover the relevant state–action space, and when small deviations from expert behaviour do not lead to severe performance degradation.
@@ 3.2 Learning from Noisy or Suboptimal Experts
Although IL relies on expert demonstrations, fully optimal and noise-free data are rarely available in practice. Human experts may be inconsistent, and computing optimal solutions can be prohibitively expensive. Consequently, most real-world IL applications rely on imperfect but readily available demonstrations. In CO, human experts are generally unsuitable due to problem complexity, and optimisation solvers are used instead. However, solving SDPs under uncertainty is computationally demanding [powell2022reinforcement], motivating the use of static approximations or simplified scenarios to generate demonstrations.
@@
Several studies in classical IL literature adopt this approach by solving static or simplified CO problems offline to obtain training targets [pham2023prediction, baty2024combinatorial]. Although such solutions are suboptimal, they are informative and significantly cheaper to compute. In many cases, these demonstrations are treated as optimal and used directly within classical IL pipelines. Nevertheless, given the $\mathcal{NP}$-hard nature of many CO problems, even static instances often require heuristics, time limits, or prescribed optimality gaps, introducing additional noise into the demonstrations. Despite its prevalence, the impact of such noise on IL performance in CO settings remains poorly understood, and systematic comparative studies are limited.
@@
Recent work has begun to address learning under noisy supervision explicitly. For example, [xu2022discriminator, kim2021demodice] propose offline IL frameworks that leverage supplementary datasets generated by suboptimal policies. [sekhari2024selective] develop an IL algorithm that reduces expert queries while achieving low regret under noisy demonstrations. [wang2023imitation] introduce a two-stage approach that denoises demonstrations using diffusion models before imitation. Additional contributions in this area include [wu2019imitation, sasaki2020behavioral, wang2021learning], as summarised in [zare2024survey].
@@
An additional challenge arises when multiple labels are available for the same state, often generated by different imperfect experts. Such situations may lead to conflicting demonstrations—for instance, when different drivers adopt distinct strategies under identical conditions. While this diversity can enrich the learning signal, it necessitates aggregation or filtering mechanisms to avoid degraded policy performance. Early work by [raykar2009supervised, raykar2010learning] addresses this issue in pre-labelled datasets. More recently, [sun2023mega] propose MEGA-DAgger, an interactive framework for learning from multiple imperfect experts that filters unsafe demonstrations and evaluates experts using scenario-specific criteria. [sekhari2024selective] further extend this line of work to settings in which different experts specialise in distinct regions of the state space. However, this line of work has not been explored when using IL for problems in operations research.
@@ 3.3 Applications of IL in SDPs
Recent applications of ML to SDPs in CO broadly fall into two categories. The first, ML-augmented optimisation, uses learning models to enhance classical optimisation algorithms. The second, known as end-to-end learning (E2EL), integrates optimisation components directly within learning architectures. This distinction follows the taxonomy of [kotary2021end]. While both IL and RL are used in these settings, we focus on work that applies IL to guide decision making within these two paradigms.
@@ 3.3.1 ML-Augmented Optimisation
ML-augmented optimisation embeds learning models within optimisation algorithms to guide or accelerate key decisions. Learning may be used to initialise algorithms, tune parameters, or repeatedly inform decisions during execution, such as branching, node selection, or column generation [bengio2021machine]. Existing approaches either predict discrete decisions directly or estimate surrogate scores that rank candidate actions and guide the search process.
@@
    Cutting Planes - Cut generation is a central technique in mixed-integer programming, aimed at tightening relaxations without excluding feasible integer solutions. Several studies use supervised learning to decide which cuts to add, including [deza2023machine, ding2020accelerating, huang2022learning, paulus2022learning].
@@
    Column Generation and Branch-and-Price - Column generation solves large-scale linear programmes by generating variables on demand. When combined with Branch-and-Bound, this yields the Branch-and-Price framework [barnhart1998branch]. Learning-based methods in this setting focus on selecting promising columns or designing effective branching strategies to improve speed. Representative works include [furian2021machine, morabit2021machine, pereira2022learning, sun2022learning, shen2022enhancing].
@@
    Large Neighbourhood Search - Large Neighbourhood Search (LNS) improves solutions by iteratively destroying and repairing partial solutions. ML has been used to generate high-quality initial solutions and to adaptively select neighbourhoods [hizir2025large, huang2023searching, liu2022learning, mossina2019multi, Nair2020SolvingMI, song2020general, sonnerat2021learning].
@@
    Search Space Reduction - Another line of work aims to directly reduce the search space by eliminating decisions unlikely to appear in an optimal solution. For example, in the travelling salesman problem, ML models can predict irrelevant arcs for pruning. Studies following this approach include [fischetti2019machine, fitzpatrick2021learning, grassia2019learning, lauri2019fine, larocca2024combining, li2025learning, spieckermann2025reduce, sun2021generalization, tahir2021improved, tayebi2022learning, vaclavik2018accelerating].
@@
    Constraint Programming - Constraint Programming (CP) solves combinatorial problems through systematic constraint propagation and search. Several works leverage supervised learning to guide CP solvers by predicting constraint satisfaction or search decisions, including [selsam2018learning, sun2024enhancing, xu2018towards].
@@
    Algorithm Configuration - ML is also used to decide whether specific algorithmic components should be applied to a given instance. Examples include predicting when to apply Dantzig–Wolfe decomposition [kruber2017learning], when to use scaling techniques [berthold2021learning], whether to apply cutting planes beyond the root node [berthold2206learning], and whether to linearise the objective in mixed-integer quadratic programmes [bonami2018learning]. In all cases, algorithm runs are used to generate labelled training data.